For the same diffusion time, the depth of penetration at 500 degree celcius and 850 degree celcius is in the ratio 1:6. The activation process for the above diffusion process is?
1
Expert's answer
2020-06-18T11:00:00-0400

The diffusion coefficient is


"D=D_0e^{-\frac{E_a}{2RT}}."

We have


"\frac{D_1}{D_2}=\frac{1}{6}=\frac{e^{-\frac{E_a}{2RT_1}}}{e^{-\frac{E_a}{2RT_2}}}."

Solve for "E_a":


"E_a=\frac{2T_1T_2R\text{ ln}(D_1/D_2)}{T_1-T_2}=36\text{ kJ/mol}."


The activation energy for the diffusion process is 36 kJ/mol.